OpenStudy (anonymous):

A cylinder has a volume of 56.52 m3 and a radius of 3 m. What is its height? (Assume = 3.14) A. 6.28 m B. 4 m C. 2 m D. 12.56 m

OpenStudy (anonymous):

\[V = \pi(r^2)h\] \[h = \frac{ V }{ \pi(r^2)}\]

OpenStudy (anonymous):

plug 56.52 for V, and 3 for r

OpenStudy (anonymous):

you should get c, i think

OpenStudy (anonymous):

ok thanks!